OpenStudy (itsmichelle29):

Help plzzzz i medal and fan The function f(x) = 4(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate?

OpenStudy (itsmichelle29):

f(x) = 4(4)^x; growth rate 400% f(x) = 4(4)^3x, growth rate 4% f(x) = 4(1.59)^3x; growth rate 59% f(x) = 4(1.59)^x; growth rate 4%

OpenStudy (campbell_st):

so is the equation \[f(x) = 4(4)^x\]

OpenStudy (itsmichelle29):

Yeah

OpenStudy (itsmichelle29):

So...... is would be f(x) = 4(4)^3x, growth rate 4% ... or no

OpenStudy (itsmichelle29):

wait so u wont change the function

OpenStudy (campbell_st):

I think you need to ask someone else... It doesn't make a lot of sense to me

OpenStudy (itsmichelle29):

okay